OpenStudy (anonymous):

How can I work out this: lim x-> 0 (xsinx)/(abs(x)) , not sure what to do with the absolute

OpenStudy (helder_edwin):

use one-sided limits

OpenStudy (anonymous):

Is there a way to do it algebraically?

OpenStudy (helder_edwin):

For instance \[ \large \lim_{x\to0+}\frac{s\sin x}{|x|}=\lim_{x\to0+}\frac{x\sin x}{x}= \lim_{x\to0+}\sin x=0 \]

OpenStudy (anonymous):

Would the squeeze theorum work here?

OpenStudy (anonymous):

oh neat, and this would be the limit for the answer? or do i have to find the one of the left as well?

OpenStudy (helder_edwin):

do the left side.

OpenStudy (anonymous):

lim x-> 0- (xsinx)/-x = -0 = 0

OpenStudy (anonymous):

does that make sense

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

oh nice thanks

OpenStudy (helder_edwin):

u r welcome